Epsom salts, a strong laxative used in veterinary medicine, is a hydrate, which means that a certain number of water molecules are included in the solid structure. The formula for Epsom salts can be written as MgSO4⋅xH2O, where x indicates the number of moles of H2O per mole of MgSO4. When 4.555 g of this hydrate is heated to 250 ∘C, all the water of hydration is lost, leaving 2.225 g of MgSO4.

What is the value of x?
Answer #1

Given,

When 4.555 g of this hydrate is heated to 250 ∘C, all the water of hydration is lost, leaving 2.225 g of MgSO4

=> In 4.555 g of this hydrate, mass of water = 4.555 - 2.225 = 2.33 g

Therefore we can say that for 2.225 g of MgSO4, 2.33 g of H2O is present in the hydrated form

Moles of 2.225 g MgSO4 = 2.225 / 120.366 = 0.0185 moles

Moles of 2.33 g water = 2.33 / 18 = 0.129 moles

=> In hydrate molecule, for 0.0185 moles of MgSO4, 0.129 moles of H2O is present

=> For 1 mole of MgSO4, number of moles of H2O present = 0.129 / 0.0185 = 7

Therefore the hydrate molecule is: MgSO4.7H2O

X = 7